OpenStudy (anonymous):

A gas occupies 200 mL at -73.0°C. To have the same gas occupy 300 mL, should the temperature be increased or decreased? What is the new temperature in Kelvins? K

OpenStudy (anonymous):

(PV)/T = (PV)/T. You can exclude the pressures since they are not mentioned. This means that V1/T1 = V2/T2. Kelvin is C + 273. Sooo, solving for T2=(V2T1)/V1. T2=(0.300L*200K)/0.200L=300K. NOTE: Volume should be in Liters, Temp in Kelvin, Pressure in atm (not used here).